### v1

# Towards a Physics Foundation Model

Abstract
Foundation models have revolutionized natural language processing through a “train once, deploy anywhere” paradigm, where a single pre-trained model adapts to countless downstream tasks without retraining. Access to a Physics Foundation Model (PFM) would be transformative—democratizing access to high-fidelity simulations, accelerating scientific discovery, and eliminating the need for specialized solver development. Yet current physics-aware machine learning approaches remain fundamentally limited to single, narrow domains and require retraining for each new system. We present the General Physics Transformer (GPhyT), trained on 1.8 TB of diverse simulation data, that demonstrates foundation model capabilities are achievable for physics. Our key insight is that transformers can learn to infer governing dynamics from context, enabling a single model to simulate fluid-solid interactions, shock waves, thermal convection, and multi-phase dynamics without being told the underlying equations. GPhyT achieves three critical breakthroughs: (1) superior performance across multiple physics domains, outperforming specialized architectures by up to 29x, (2) zero-shot generalization to entirely unseen physical systems through in-context learning, and (3) stable long-term predictions through 50-timestep rollouts. By establishing that a single model can learn generalizable physical principles from data alone, this work opens the path toward a universal PFM that could transform computational science and engineering.
1 Introduction
Over the last few years, massive accumulation of data and large-scale GPU computing have led to extraordinary advancements in language model (LLMs) capabilities (Devlin et al., 2019; Radford et al., 2018; 2019; Raffel et al., 2020; Anil et al., 2023). These frontier foundation models, often exceeding 100 billion parameters, have established a “train once, deploy anywhere” paradigm. They generalize to unseen domains and can be prompted to perform diverse tasks—from coding to creative writing—without task-specific finetuning (Brown et al., 2020; Minaee et al., 2025), exhibiting emergent abilities not explicitly programmed (Wei et al., 2022).
We envision a similar paradigm for physics-aware machine learning (PAML), where a single foundation model can simulate a wide range of physical systems, boundary conditions, and initial states. This allows end-users to employ the model for their individual use case without the need for extensive retraining or fine-tuning. However, current state-of-the-art physics models, such as physics-informed neural networks (PINNs) (Raissi et al., 2019), neural operators (Kovachki et al., 2023; Lu et al., 2021), and physics-aware recurrent convolutions (PARC) (Nguyen et al., 2023a; 2024), are fundamentally limited to solving a single, narrowly scoped physical system: While they excel at their specific task, they cannot generalize to new physics or boundary conditions without new or additional training. Even recent multi-physics approaches largely rely on fine-tuning or meta-learning, which still demands new data and training for each application (Penwarden et al., 2023; Cho et al., 2023; McCabe et al., 2024).
The primary barrier to a truly “train once, deploy anywhere” Physics Foundation Model (PFM) is the immense diversity of physical phenomena coupled with highly expensive and limited data. A single model intended as a surrogate for computational fluid dynamics, for example, must reconcile the micrometers and milliseconds of microfluidics with the kilometers and hours of weather forecasting. Critically, the same initial state can evolve into vastly different outcomes depending on the governing physical laws. A PFM must therefore either be explicitly provided with a complete system description (e.g. scale, boundary conditions, material properties, governing equations) or infer the dynamics from the given input data itself. As the variety of problems scales, the former becomes impractical and defeats the purpose of a truly general model.
Here, we propose that the path towards PFMs lies in emulating the in-context learning abilities of LLMs (Brown et al., 2020; Agarwal et al., 2024). Instead of being explicitly told the governing equations, a model should infer the underlying dynamics from a "prompt" consisting of a short sequence of prior states. Such a model could adapt its predictions on the fly, enabling a single, unified architecture to tackle a wide array of physical scenarios. This paradigm shift presents fundamental research challenges that we address through the General Physics Transformer (GPhyT) trained on a diverse 1.8 TB corpus of simulation data. With this model, we investigate three critical questions:
  1. Q1:
    Can a single, large-scale transformer effectively model a wide range of disparate physical systems (e.g., incompressible flow, shock waves, convection) without any explicit, physics-describing features? (Section 4.1)
  2. Q2:
    Can this foundation model perform zero-shot generalization to new, unseen physical conditions (e.g., new boundary conditions, entirely new physics) by inferring the dynamics from the input alone? (Section 4.2)
  3. Q3:
    Can GPhyT maintain physical consistency and stability during extended autoregressive rollouts, a characteristic crucial for real-world application? (Section 4.3)
Our results demonstrate that GPhyT not only outperforms specialized architectures on seen tasks but also successfully generalizes to out-of-distribution problems, including producing physically plausible predictions for phenomenon absent from its training data. This work represents a critical step towards creating a "universal physics engine" that could democratize access to high-fidelity simulations and accelerate scientific discovery across disciplines.
2 Related work
Neural Surrogates for Physical Systems
Machine learning has emerged as a powerful tool to accelerate the simulation of complex physical systems, which are governed by partial differential equations (PDEs) that lack analytical solutions. The dominant paradigms in this domain are Physics-Informed Neural Networks (PINNs), Neural Operators (NOs) and their combination Physics-informed Neural Operators. PINNs embed the governing PDEs directly into the training process as a soft constraint in the loss function, which enhances data efficiency and physical consistency (Raissi et al., 2019; Karniadakis et al., 2021). This approach has been successfully applied across numerous scientific fields (Faroughi et al., 2024). Neural Operators, in contrast, learn the solution operator mapping from the PDE parameters to the solution space, making them discretization-invariant (Kovachki et al., 2023). Prominent examples include Fourier Neural Operators (FNOs), which perform convolutions in the frequency domain (Li et al., 2020), and DeepONets (Lu et al., 2021). Moreover, combinations of operators with physics-informed loss functions can reduce the data requirements of neural operators (Goswami et al., 2023; Li et al., 2023).
Despite their success, PINNs and NOs are fundamentally specialized solvers. They are typically designed and trained for a single, well-defined physical system and struggle to generalize to new governing equations, boundary conditions, or complex multi-physics phenomena without transfer learning (Goswami et al., 2022; 2020) or full retraining. This inherent specialization prevents them from serving as true "foundational" models in the way that large language models (LLMs) do for natural language tasks.
Towards Foundational Models for Science
The concept of a large-scale foundation model pretrained on extensive, diverse data has begun to permeate scientific disciplines. This has led to two distinct categories of models. The first involves language-based models fine-tuned on scientific corpora, such as AstroLLaMA for astronomy (Nguyen et al., 2023b) or specialized models for interpreting medical records (Jiang et al., 2023). The second category comprises models that operate directly on quantitative scientific data such as velocity or temperature fields. Notable examples are models for molecular structures (Chithrananda et al., 2020), climate forecasts (Nguyen et al., 2023c), or aquatic science (Yu et al., 2025). Regardless of methodology, any foundation model must either be capable generalizing to unseen data or finetuned for new tasks (Choi et al., 2025).
In physics, the pursuit of foundation models has largely focused on enhancing the generalization of neural surrogates. Researchers have explored meta-learning (Penwarden et al., 2023) and transfer learning (Subramanian et al., 2023; Goswami et al., 2022) to adapt pretrained models to new PDE systems with fewer data samples. Recently, multi-tasks models trained on multiple physical systems were explored McCabe et al. (2024). However, while both groups demonstrated superior accuracy compared to single-physics models, both opted for finetuning to unseen tasks. These efforts represent important progress, but they still fall short of the “train once, deploy anywhere” paradigm that we envision.
Transformers for Spatiotemporal Modeling
The architectural backbone of most modern foundation models is the Transformer (Vaswani et al., 2017), whose self-attention mechanism has proven exceptionally effective at capturing long-range dependencies in sequential data. Originally developed for language, this architecture was successfully adapted for computer vision in the Vision Transformer (ViT) (Dosovitskiy et al., 2020). By treating an image as a sequence of patches, ViTs achieved state-of-the-art performance with sufficient data provided (Khan et al., 2023). This concept was further extended to video by creating spatiotemporal "tubelet" tokens (Arnab et al., 2021), enabling transformers to model dynamic visual data. The power of transformers also extends to generative tasks. Using vector quantization, auto-regressive transformer models can operate on a discrete latent space of visual tokens (Esser et al., 2021; Chang et al., 2022; Ramesh et al., 2021).
3 General Physics Transformer
3.1 Architecture
Due to data scarcity, today’s physics models must incorporate inductive biases for optimal performance. However, the diversity of multiple physical systems restricts such choices. The General Physics Transformer (GPhyT) is designed as a hybrid model that integrates a deep learning component within a classic numerical methods framework. As illustrated in Figure 1a, the core of our architecture is a Transformer-based neural differentiator that learns the temporal dynamics of a system, coupled with a standard numerical integrator that extrapolates the system’s future state. This approach, inspired by Neural ODEs (Chen et al., 2018) and previous work of Nguyen et al. (2024), allows the model to predict the evolution of diverse physical systems governed by partial differential equations (PDEs).
Neural differentiator
The neural differentiator (blue dashed box) models the partial derivative ($\frac{\partial X}{\partial t}$) of the physical state with respect to time. $X$ is composed for multiple physical fields (channels), such as pressure, temperature, and velocity. To allow for in-context learning, the differentiator receives multiple time snapshots ($X_{t_{i}-n},...,X_{t_{i}}$) of the physical state. The sample is then tokenized by a single linear transformation across spatial and temporal dimensions, yielding non-overlapping spatiotemporal (tubelet (Arnab et al., 2021)) patches. The size of these patches control the number of spatial and temporal pixels encoded in each token. Absolute positional encodings are added to the patches. The spatiotemporal transformer consists of multiple transformer layers with layer norms and attention across all time and space dimensions, illustrated in Figure 1b. We chose this unified attention mechanism over more computationally efficient factorized approaches to ensure maximum expressivity, allowing the model to capture complex, non-separable phenomena like turbulence and shockwave interactions. Finally, a linear transformation (detokenizer) reverts the spatiotemporal patches into the input space.
To provide the model with explicit local information, we compute the first-order spatial ($dx,dy$) and temporal ($dt$) derivatives of the input fields using central differences. These computed derivatives are concatenated with the original fields along the channel dimension, enriching the input for the neural differentiator. This technique is particularly effective for resolving phenomena with sharp gradients (Cheng et al., 2024).
Figure 1: (a) General architecture of GPhyT. A 4D-stack of physical quantities (time, height, width, fields) serves as input $X$. The numerically computed derivatives of each field are concatenated to the input. The differentiator (linear tokenizer, spatiotemporal transformer, linear detokenizer) provides the partial derivative of $X$ wrt. time. Finally, a numerical integrator computes the next timestep of each field given $\frac{\partial X}{\partial t}$ and $X$. (b) Architecture of a single transformer layer, consisting of layer norms (LN), spatiotemporal attention, and multilayer perceptron (MLP).
Numerical Integrator
With the learned time derivative, we can predict the next state of the system, $X_{t_{i+1}}$, using a numerical integration step. The general form of the integration is:
$X_{t_{i+1}}=f\left(X_{t_{i}},\frac{\partial X}{\partial t}\bigg{|}_{t_{i}},\Delta t\right)$ (1)
In this study, we choose the first-order Forward Euler method. An ablation study in Appendix 6.4 showed that higher-order integrators like Runge–Kutta 4 provided no significant accuracy benefit for our framework, making Forward Euler the most computationally efficient choice.
3.2 Datasets
To train a model capable of learning general physical principles, we curated a large and diverse corpus of simulation data, comprising eight distinct datasets listed in Table 1. The combined dataset contains over 2.4 million simulation snapshots, totaling 1.8 TB of data. Our data is sourced from both the publicly available "The Well" benchmark (Ohana et al., 2024) and our own custom simulations, described in detail in Appendix 6.6.
The three datasets from The Well cover a range of fundamental physics, including incompressible (Shear flow) and compressible (Euler) fluid dynamics, and thermal convection (Rayleigh–Bénard). However, these systems largely lack the solid boundaries and complex geometries prevalent in engineering applications. To address this, we generated four additional datasets featuring flows around rigid obstacles, Rayleigh–Bénard with additional obstacles, heat exchange with solid elements (Thermal Flow), and multi-phase dynamics in porous media. These additions introduce critical physical behaviors, such as boundary layer formation, vortex shedding, pressure-driven instabilities, as well as varying physical scales, significantly expanding the diversity of the training data.
Table 1: Dataset overview. Each trajectory has different initial conditions / randomized geometry and contains the given number of timesteps. Unique samples are all possible combinations of 4 input and 1 output snapshots sampled for time-increments of 1-8 and random axis flips.
Dataset Trajectories Timesteps Snapshots Unique samples Origin
Shear flow 1120 200 224,000 6,522,880 The Well
Rayleigh–Bénard 1750 200 350,000 10,192,000 The Well
Euler 5000 100 500,000 13,120,000 The Well
Obstacle flow 1266 481 608,946 18,756,856 Own
Thermal flow 354 481 170,274 5,244,864 Own
Rayleigh–Bénard 2 228 1001 228,228 7,171,968 Own
Twophase flow 816 401 327,216 10,000,896 Own
Sum 2,408,664 71,009,464
A core objective of this work is to train a model that can generalize by inferring the underlying physics from context. To facilitate this, we implemented two crucial data augmentation strategies:
  • Variable Time Increments: Each simulation trajectory is sub-sampled using multiple time-step increments ($\Delta t$). This forces the model to learn dynamics that are invariant to the sampling frequency. For any given input, the model must infer the temporal scale from the dynamics presented in the prompt, as a single time step could represent milliseconds in one context and minutes in another.
  • Per-Dataset Normalization: The physical phenomena in our corpus span vastly different scales, from micrometer-sized pores in two-phase flow to large-scale convective cells. To handle this, we normalize each dataset independently. This preserves the relative physical quantities within a single simulation while compelling the model to infer the absolute magnitudes and spatial scales of a new system purely from the context provided by the input snapshots.
By training on this varied data, GPhyT is explicitly pushed to develop in-context learning abilities, rather than memorizing the characteristics of a single, fixed physical system.
4 Results
4.1 Multi-physics learning
Figure 2: Average and median mean squared error (MSE) across all test sets for each model (lower is better). Number of parameters (in million) are given below model names.
To answer our first research question Q1—whether a single model can effectively learn to represent numerous, disparate physical systems—we evaluated GPhyT’s single-step prediction accuracy across our entire multi-physics test set. We benchmarked against two established architectures: the Fourier Neural Operator (FNO) (Kovachki et al., 2023), and a standard UNet, a widely-adopted baseline for spatiotemporal prediction tasks. All models were trained identically on 4 initial frames to predict the subsequent frame, ensuring a fair comparison.
Figure 2 presents the average and median mean squared error (MSE) across all test sets. GPhyT demonstrates substantial performance gains in average MSE and achievs a 5× reduction in median MSE compared to UNet and a 29× reduction compared to FNO at equivalent model sizes (size M). Detailed per-dataset losses are provided in SI section 6.2. Remarkably, even our smallest variant (GPhyT) with only 10% of the parameters outperforms the full-sized FNO and UNet when comparing median MSE. This efficiency gain suggests that the spatiotemporal attention mechanism captures cross-scale physical interactions more effectively than traditional convolutional or neural operator approaches. Furthermore, GPhyT exhibits favorable scaling properties with performance gains from the S to XL variants, indicating potential for further improvements with increased model capacity.
Figure 3: Next timestep prediction (velocity magnitude) of GPhyT, Unet, and FNO in their M-variant against ground truth (GT). Top row: Incompressible flow with periodic boundary conditions and $\Delta t=8$. 2nd row: Shear flow ($\Delta t=4$). 3rd row: Euler shock waves with periodic boundary conditions ($\Delta t=4$). 4th row: Rayleigh–Bénard convection ($\Delta t=4$). The results can be best viewed on a high-definition digital monitor.
Figure 3 provides qualitative comparisons across representative physical systems, categorized by complexity. For smooth systems (top two rows: obstacle flow and shear flow), GPhyT and UNet precisely capture fine-scale vortical structures and boundary layer dynamics, while the FNO struggles with spatial localization. For more challenging systems featuring discontinuities and chaotic dynamics (bottom two rows: Euler shocks and Rayleigh–Bénard convection), the performance gap widens dramatically. GPhyT maintains sharp shock fronts and accurately tracks convective plumes, while the FNO fails to resolve these critical features, producing overly smoothed predictions that miss essential physical phenomena. The UNet produces reasonable approximations but exhibits some diffusion of sharp features. These results suggest that the GPhyT architecture, with its ability to dynamically attend to relevant spatiotemporal features, is inherently better suited as a Physics Foundation Model than FNO and UNet.
4.2 Zero-shot in-context learning
The defining characteristic of a true foundation model is its ability to adapt to new tasks without additional training—a capability that fundamentally distinguishes them from traditional specialized models. In language models, this emerges through in-context learning, where models leverage prompts to perform tasks never explicitly seen during training (Brown et al., 2020). To investigate our research question Q2, whether GPhyT exhibits similar emergent capabilities for physics, we designed increasingly challenging generalization experiments: First, we evaluated the model on systems with modified boundary conditions completely absent from the training data. Second, we pushed the boundaries further by presenting entirely novel physical phenomena, including supersonic flows and turbulent radiative layers never encountered during training. These experiments probe whether the model has learned transferable physical principles rather than memorizing dataset-specific patterns.
Table 2: Average MSE for in-context learning tasks. The model is evaluated on systems with new boundary conditions and entirely new physics, with known systems provided as a baseline.
Dataset Variant Zero-shot MSE
Obstacle flow
   Symmetric Known $2.97\times 10^{-4}$
   Periodic Known $5.32\times 10^{-4}$
   Open New $8.01\times 10^{-3}$
Euler
   Periodic Known $8.92\times 10^{-3}$
   Open New $9.09\times 10^{-3}$
Turbulent layer New $2.05\times 10^{-1}$
Supersonic New $1.40\times 10^{-1}$
The quantitative results are summarized in Table 2, with qualitative examples shown in Figure 4 and Section 6.8 in the Appendix. For incompressible flow around obstacles, introducing a new "open" boundary condition, where mass is not conserved, results in an average MSE nearly identical to the MSE achieved on the "known" symmetric system. Even more striking is the result for the chaotic Euler system, where the open boundary variant achieves the same accuracy compared to the known periodic system. Thus, GPhyT successfully infers new boundary conditions from the prompt without the need for finetuning.
The most challenging tests are two completely new physical systems: Supersonic flow around an obstacle and a turbulent radiative layer. Here, the MSE is significantly higher than in the known systems, emphasizing the significant challenge. Yet, Figure 4 shows that GPhyT still captures the essential dynamics, correctly forming a bow shock and the overall flow field for the supersonic flow. Similarly, the general turbulent features are present in the turbulent radiative layer. Naturally, finetuning the model on the new systems as done in previous studies McCabe et al. (2024); Subramanian et al. (2023) will yield even more accurate predictions, however, the ability to extrapolate and produce a physically plausible result for entirely new physics, even with reduced accuracy, is a powerful demonstration of emergent generalization.
Figure 4: Qualitative results for in-context learning. GPhyT accurately predicts the next state for an incompressible flow and Euler flow with unseen open boundary conditions (1st and 2nd column). For the completely new physics of supersonic flow (3rd column), it successfully captures the formation of a bow shock wave, demonstrating strong zero-shot generalization. For the turbulent radiative layer (4th column), general turbulent features are present while finegrained details are missing. The results can be best viewed on a high-definition digital monitor.
4.3 Long-range Prediction
As stated in research question Q3, the true utility of any physics surrogate model is measured by its ability to maintain stability and accuracy over extended temporal horizons. This task is exceptionally challenging, as it requires the model to generate a full trajectory from an initial state, with prediction errors from each step accumulating over time. It is a critical test of a model’s physical consistency and a common failure point that is often omitted in literature.
Figure 5: (a) Autoregressive rollout prediction for velocity magnitude of GPhyT-XL on the Euler shockwaves with periodic boundary conditions. (b) Average mean squared error (MSE) over 50 time steps for all known physical systems. (c) Average MSE over 50 time steps for the novel unseen physical systems. The results can be best viewed on a high-definition digital monitor.
Figure 5a) illustrates a 50-timestep autoregressive rollout for an Euler shockwave. In the initial phase (t=16), GPhyT-XL maintains high fidelity, accurately capturing the intricate vortex structures and shock fronts of the ground truth. As the rollout progresses, error accumulation becomes visible as a loss of high-frequency detail (t=32). However, the model successfully preserves the global dynamics and physical plausibility of the flow; the large-scale structures and boundary interactions remain stable and coherent. The quantitative evaluation across the known physical systems is shown in Figure 5b). For most physics, GPhyT’s exhibit near-linear error accumulation over 25 timesteps. Exceptions are twophase flow, where changes can occur extremely abrupt, and shear flow, where a chaotic initial system converges to a quasi-static solution.
Extension to out-of-distribution systems (Figure 5c) provides further insight into the model’s generalization mechanisms. Systems with modified boundary conditions maintain error trajectories comparable to their training counterparts. Specifically, Euler with open boundaries tracks the periodic case almost identically, suggesting the model has disentangled bulk dynamics from boundary effects. For entirely novel physics error growth rates are higher, highlighting the need for even more accurate and stable physics foundation models.
5 Conclusion
We have demonstrated that a single transformer-based model can effectively learn and predict the dynamics of diverse physical systems without explicit physics-specific features, marking a significant step toward true Physics Foundation Models. GPhyT not only outperforms specialized architectures on known physics by up to an order of magnitude but, more importantly, exhibits emergent in-context learning capabilities—inferring new boundary conditions and even entirely novel physical phenomena from input prompts alone. This “train once, deploy anywhere” capability, previously exclusive to language models, opens new possibilities for physics simulation. Our hybrid architecture, combining a transformer-based neural differentiator with numerical integration, proves that the attention mechanism can capture complex spatiotemporal dependencies across vastly different scales. The model’s ability to maintain physical consistency through 50-timestep rollouts, while not yet matching numerical solvers, demonstrates that learned representations can encode generalizable physical principles rather than merely memorizing dataset-specific patterns. The path toward a comprehensive Physics Foundation Model requires addressing current limitations: extending to 3D systems, incorporating diverse physical domains beyond fluid dynamics, and achieving variable-resolution capabilities. Most critically, improving long-term stability will be essential for practical engineering applications. Nevertheless, GPhyT establishes that the foundation model paradigm, i.e. a single pre-trained model adapting to novel tasks through context alone, is achievable for physics. As we scale both model capacity and training data diversity, we anticipate further emergent capabilities that could fundamentally transform how we approach computational physics, making high-fidelity simulations accessible to researchers and engineers without the traditional barriers of specialized solver development or extensive computational resources.
Acknowledgments
F.W acknowledges financial support from the German Federal Ministry of Research, Technology and Space (BMFTR project “PrometH2eus”, FKZ 03HY105A). M.W. acknowledges DFG funding through the Gottfried Wilhelm Leibniz Award 2019 (WE 4678/12-1). S.B. acknowledges financial support from the National Science Foundation (Award No. DMREF-2203580). The authors acknowledge Research Computing at the University of Virginia for providing computational resources that have contributed to the results reported within this paper. Additional computational resources were provided by the German AI Service Center WestAI. The authors acknowledge the use of LLMs for improved grammar and wording in the manuscript. Furthermore, they acknowledge the use of coding agents (e.g. Claude code / Cursor) to generate boilerplate code and to draft unit tests. The views and conclusions of this work are those of the authors only.
Reproducibility statement
All training, evaluation and plotting code will be published. Furthermore, model checkpoints will be released on Hugging Face. Finally, the datasets which are not available on The Well will be avaiable to download.